# SOPHY: Generating Simulation-Ready Objects with PHYsical Materials

### v1

@@ Results.
SOPHY outperforms the other methods, demonstrating positive z-scores for alignment with real-world physics based on VideoPhy. In terms of image similarity, our method is better aligned with the ground-truth. We also show qualitative results for a few sample frames for test scenes in Figure 5. We observe that DG4D struggles to generate noticeable deformations, even for soft objects e.g., plants, due to its reliance on Stable Video Diffusion (SVD) [3], which is relatively hard to tune for deformation. The baseline model generates object geometry independently of physical materials. Thus, it often creates physically implausible objects or predicts inappropriate materials for the generated shapes.  SOPHY produces more realistic dynamics aligning closely with the given conditions.
@@ 5.3 Text-to-4D
We also evaluate  SOPHY for text-to-4D. The evaluation is the same as before with the only difference being the input condition (text instead of an image). We test on prompts, describing the object category, part tags, fine-grained material categories for our test objects, following the prompts of Section 3.1. We also include the test scenario in the prompt.
@@
We compare our method with the text-to-4D method of STAG4D [60] that generates deforming 3DGS driven by text prompts. We note that our method outputs 4D videos about $10$x faster than STAG4D based on the same GPU and splat renderer. We also compare with our baseline. For evaluation metrics, we use the z-normalized VideoPhy scores and CLIP similarity for measuring alignment of the rendered frames with the input text prompt averaged over all frames and test scenes.
@@ Results.
SOPHY achieves the highest performance in VideoPhy Score, demonstrating a large positive score in contrast to STAG4D’s negative score. The CLIP score is comparable for all methods demonstrating similar alignment with text. We also present a visual comparison for the generated dynamics in Figure 6. We observe that STAG4D produces unrealistic object behavior, e.g., in Figure 6(b), the size of the handbag fluctuates over time. Compared to the baseline, our method produces more geometrically coherent and physically plausible results. These findings highlight SOPHY’s ability to generate simulation-ready 3D objects for diverse, physically realistic 4D content.
6 Discussion
We have presented a new generative model of 3D objects incorporating geometry, color, and physical material properties. Our experiments demonstrated significant benefits of the approach, including generating physically plausible 4D videos from images and text. Interesting avenues for future work include extending our generative model for synthesizing whole scenes and generating other phenomena e.g., fluids or gases. More accurate simulators, e.g., based on finite elements, could be used instead since our model generates the full objects’ volume. Our annotated dataset is only a first step towards more physics-aware 3D datasets – enlarging it with more objects and material properties would be a fruitful direction.
Acknowledgements.
@@
Supplementary Material
7 Full Prompts
@@
In this section, we provide the detailed prompts we used in the VLM-guided material annotation stage described in Section 3.1.
7.1 Fine-grained Material Types
@@
part_name’: most_suitable_material_type’,
@@
7.2 Material Models and Parameters
@@ 3. Drucker-Prager plasticity (Youngs modulus, Poissons ratio, friction angle);
For each part, you need to specify **both the elastic and the plastic** material model and the material parameters.
Please provide your answer in the following JSON format (for Youngs modulus and yield stress, the unit is Pa.):
@@
"...": ..., // other parameters, e.g., yield stress ("sigma_y"), friction angle ("phi")
@@
7.3 Update with Expert Feedback
In this case, we send three messages with the role set by user, assistant, user in succession. The first user prompt is the same as the prompt for obtaining the initial material properties. The assistant prompt is the output of the original query from the VLM. The second user prompt is as follows.
@@ 7.3 Update with Expert Feedback
8 Dataset Details
In Table 4, we present the detailed composition of the proposed dataset, which covers $12$ common categories with detailed part-level material annotations for $3,004$ shapes. After the VLM-guided material annotation, we obtained $8$, $7$, $10$ types of fine-grained materials for “fabric”, “leather”, and “plastic” respectively [^2] We did not obtain parts labeled with “bonded leather” for the “leather” type, and parts labeled with “flexible polyvinyl chloride”, “polystyrene” for the “plastic” type from the VLM.. The occurrence of parts labeled with these fine-grained materials is illustrated in Figure 7. We also present the statistics of material behavior types in the proposed dataset in Figure 8. Specifically, M0 represents pure elastic materials using neo-Hookean elasticity and identity plasticity. M1 represents plastic materials with softening implemented with neo-Hookean elasticity and von Mises plasticity with damage. M2 is for plastic material without softening, which is supported by neo-Hookean elasticity and von Mises plasticity. Finally, M3 stands for granular material with StVK elasticity and Drucker-Prager plasticity as the material models.
Our dataset offers detailed material properties, which serve as a first step for advancements in simulation-driven learning, physics-informed generative modeling, and interactive virtual environments. We hope that our dataset can expand the scope of 3D perception research and contribute to developing effective algorithms and techniques that improve our understanding of the physical world.
Figure 7: Statistics of fine-grained material types in our dataset.
Figure 8: Statistics of material models in our dataset.
Train 75 418 898 27 18 27 139 71 383 278 37 91 2462
Valid 10 26 52 5 2 5 21 11 21 18 6 3 180
Test 24 48 106 8 7 10 39 18 46 31 11 14 362
Table 4: Data composition of our proposed dataset.
Figure 9: Examples from our proposed dataset.
9 Implementation Details
9.1 Denoiser and Conditional Signals
In Figure 10, we present the denoising network used in our method. We alternatively use a self-attention (SA) layer and a cross-attention (CA) layer for image- and text-conditioned generation. The cross attention is used to inject the conditional signal $\boldsymbol{c}$ into the latent representation. In the image-conditioned experiment, we use 10 rendered images from randomly sampled viewpoints with a size of $256\times 256$ as training data. In the text-conditioned experiment, each object is accompanied by 5 text descriptions. 4 of these descriptions are automatically generated using metadata from the 3DCoMPaT200 [1], as shown in the code snippet below. The fifth description is created by querying ChatGPT-4o [37] for a text description of the object based on two rendered views.
1# Conditional Signals for Text
2# 1. with color, coarse material type, and part label
3text_1 = f"A {shape_name} made of"
5 mat_color = MAT_COLORS[mat].lower()
6 if i != len(part_descriptions):
7 text_1 += f" a {mat_color} {mat} {part},"
8 else:
9 text_1 += f" and a {mat_color} {mat} {part}."
10
11# 2. with fine material type and part label
12text_2 = f"A {shape_name} made of"
14 if mat_fine is None:
15 mat_name = mat
16 else:
17 mat_name = mat_fine
18 if i != len(part_descriptions):
19 text_2 += f" a {mat_name} {part},"
20 else:
21 text_2 += f" and a {mat_name} {part}."
22
23# 3. with part label
24text_3 = f"A {shape_name} composed of"
26 if i != len(part_descriptions):
27 text_3 += f" a {part_name},"
28 else:
29 text_3 += f" and a {part_name}."
30
31# 4. with category label
32text_4 = f"A 3D shape of {shape_name}."
Figure 10: Denoiser architecture. We train the denoising network in the latent space (refer to Equation 3) and adopt alternative self attention (SA) and cross attention (CA) to learn denoised representations. The cross attention is for incorporating conditional signal ($\boldsymbol{c}$) like images or text prompts.
9.2 Data Preprocessing
For object autoencoding and generation, we mainly follow the preprocessing pipeline of 3DShape2VecSet [62] to obtain per-point information. First, each 3D shape is converted into a watertight mesh using ManifoldPlus [16] and normalized to its bounding box. From this, we sample a dense surface point cloud of $150,000$ points. To train the network, we randomly sample $300,000$ points in 3D space—each annotated with occupancy, color, and material properties—along with an additional $300,000$ points from the near-surface region with the same attributes. We apply the frequency embedding [35, 62] to each point’s position and color before forwarding it to the encoder. We note that per-point color is assigned using a 1-NN approach, where each point inherits the color of its nearest surface point. Material properties are determined through a label propagation process. 3DCoMPaT200 [1] provides surface points with part labels for each object. To extend these labels to volumetric points, we apply a 5-NN algorithm, assigning labels via majority vote. Once part labels are obtained, we map corresponding material attributes based on shape part annotations in our material-augmented dataset.
9.3 Training Details
For object autoencoding, we use a surface point cloud of $2,048$ points as input to the autoencoder. At each iteration, we sample $1,024$ query points from the bounding sphere and another $1,024$ from the near-surface region for attribute prediction. Following [61], we ensure an equal distribution of query points between occupied and non-occupied regions. We set the weight coefficients described in Equation 8 as $\lambda_{c}=\lambda_{\nu}=\lambda_{\phi}=\lambda_{M}=0.1$, $\lambda_{E}=\lambda_{\sigma}=0.01$, and $\lambda_{r}=0.001$. The autoencoder is trained with a batch size of $256$ for $500$ epochs, using a learning rate that linearly increases to $10^{-4}$ over the first $25$ epochs before gradually decaying to $10^{-5}$ following a cosine schedule.
For object generation, we train our diffusion models with a batch size of $256$ for $2,000$ epochs. The learning rate is linearly increased to $2\times 10^{-4}$ over the first $200$ epochs and then gradually decays to $10^{-6}$ using a cosine schedule. We adopt the default hyperparameter settings of EDM [19]. During sampling, the final latent codes are obtained using only 18 denoising steps.

### v2

@@ Results.
In Section 5, SOPHY outperforms the other methods, demonstrating the highest z-score for alignment with real-world physics based on VideoPhy. For CLIP scores, our method is better aligned with the conditioned signals. We also show qualitative results for a few samples for test scenes in Figure 5. We observe that DG4D struggles to generate physically plausible animations, e.g., the headband size fluctuates over time, as it overlooks physical constraints. Both baselines, “B. Dec.” and “B. Perc.”, generate geometries and physical materials independently. Thus, they often create physically implausible objects (e.g., strange chairs in Figure 5(b)) or predict inappropriate materials for shapes (e.g., too stiff materials for a bag that looked like it was made of fabric in Figure 5(c)). In contrast, SOPHY produces more realistic dynamics, aligning closely with the given images.
Figure 6: Qualitative comparison in the text-to-4D setting. The text prompts used as input conditions are shown below. Note that the ground plane is not shown here as well.
@@ 5.3 Text-to-4D
We also evaluate SOPHY for text-to-4D generation. The evaluation is the same as before with the only difference being the input condition (text instead of an image). We test on prompts, describing the object category, part tags, fine-grained material categories for our test objects, following the prompts of Section 3.1. We also include the test scenario in the prompt.
@@
We compare our method with STAG4D [84], which generates deforming 3DGS driven by text prompts. We also consider the two baselines described in the previous section here. The evaluation metrics are the same as those used in the image-to-4D setting.
@@ Results.
In Section 5, it is observed that SOPHY achieves the highest performance in VideoPhy and CLIP scores, demonstrating better alignment with real-world physics and given text prompts. We also present a visual comparison for the generated dynamics in Figure 6. We observe that STAG4D fails to produce noticeable deformations, even for soft objects like teddy bears. Since the two baselines exclude material properties from the generation process, they fail to produce appropriate physical attributes aligned with the text prompts. For instance, in Figure 6(a), “B. Dec.” generates a very stiff plant while “B. Perc.” creates a soft metal container. Compared to them, our method produces more geometrically coherent and physically plausible results.
6 Conclusion
We presented a new generative model of 3D objects incorporating geometry, color, and physical material properties. Our experiments demonstrated significant benefits of the approach, including generating physically plausible 4D videos from images and text. Interesting avenues for future work include extending our generative model for synthesizing whole scenes and generating other phenomena e.g., fluids [36, 48, 69] or gases [17]. More accurate simulators, e.g., based on finite elements [62, 4, 68], could be used instead since our model generates the full objects’ volume. Our annotated dataset is only a first step towards more physics-aware 3D datasets – enlarging it with more objects and material properties would be a fruitful direction.
Acknowledgments
@@
Appendix A Additional Analysis
A.1 Conditioning on images from other datasets
Here, we investigate how well SOPHY generalizes to conditioned images sourced from different datasets. We specifically select the Objaverse [16], a widely used dataset in previous 3D generation research, to obtain 2D renderings of several 3D objects that serve as input conditional signals for our method. After completing the physics-aware 3D generation, we simulate the generated assets using the pipeline described in Section 3.2 of the manuscript, finally obtaining dynamic videos. We present the results in Figure 7. Even though our method was not specifically trained on this dataset, it can still generate 3D assets with plausible material properties and physical simulations.
Figure 7: Generated results by conditioning our method on images from Objaverse [16] shown on the left.
Figure 8: Changing parts or object categories in text prompts shown on the left. On the right, we show different samples of objects generated by our method based on the edited prompts.
A.2 Conditioning on text prompt variations
In this subsection, we study whether SOPHY can generate different 3D objects aligned well with the prompts based on edited text conditions. Specifically, we qualitatively demonstrate changes in the generated results by changing object parts or categories in the input prompt. The results are shown in Figure 8. Starting with the text shown in Figure 8(a), we add a description of “a polyester pillow” in (b). The resulting generations accurately reflect this change in the descriptions. Furthermore, we alter the object category from “sofa” in (a) to “chair” in (c). As shown in Figure 8(c), SOPHY can generate similar shapes for the unedited parts in the prompt, such as “cushion” and “armrest”, while the rest of the shape changes to reflect the edited category.
Appendix B Implementation Details
Figure 9: Denoiser architecture. We adopt self attention (SA) and cross attention (CA) blocks to learn denoised representations starting from latents (see Equation 3). The cross attention incorporates conditions $\boldsymbol{c}$ (images/prompts).
B.1 Denoiser Architecture
In Figure 9, we present the architecture of the denoising network used in our method. We alternate self-attention (SA) and cross-attention (CA) layers. The cross attention is used to inject the condition signal $\boldsymbol{c}$ (image or text prompt) into the latent representation.
B.2 Conditioning Signals for training
In the image-conditioned experiment, we use $10$ rendered images from randomly sampled viewpoints with a size of $256\times 256$ for training. In the text-conditioned experiment, each object is accompanied by $5$ text descriptions. $4$ of these descriptions are automatically generated using metadata from the 3DCoMPaT200 [1], as shown in the code snippet below. The fifth description is created by querying ChatGPT-4o [50] for a text description of the object based on two rendered views.
1# Conditional Signals for Text
2# 1. with color, coarse material type, and part label
3text_1 = f"A {shape_name} made of"
5 mat_color = MAT_COLORS[mat].lower()
6 if i != len(part_descriptions):
7 text_1 += f" a {mat_color} {mat} {part},"
8 else:
9 text_1 += f" and a {mat_color} {mat} {part}."
10
11# 2. with fine material type and part label
12text_2 = f"A {shape_name} made of"
14 if mat_fine is None:
15 mat_name = mat
16 else:
17 mat_name = mat_fine
18 if i != len(part_descriptions):
19 text_2 += f" a {mat_name} {part},"
20 else:
21 text_2 += f" and a {mat_name} {part}."
22
23# 3. with part label
24text_3 = f"A {shape_name} composed of"
26 if i != len(part_descriptions):
27 text_3 += f" a {part_name},"
28 else:
29 text_3 += f" and a {part_name}."
30
31# 4. with category label
32text_4 = f"A 3D shape of {shape_name}."
B.3 Data Preprocessing
For object autoencoding and generation, we mainly follow the preprocessing pipeline of 3DShape2VecSet [85] to obtain point-based shape representations. First, each 3D shape is converted into a watertight mesh using ManifoldPlus [23] and is normalized to its bounding box. Then we sample a dense surface point cloud of 150K points. To train the network, we randomly sample 300K points in 3D space—each annotated with occupancy, color, and material properties—along with an additional set of 300K points from the near-surface region with the same attributes. We apply the frequency embedding [47, 85] to each point’s position and color before forwarding it to the encoder. We note that color for off-surface points is assigned using a $1$-NN approach, where each point inherits the color of its nearest surface point.
Material properties are determined through a label propagation process. 3DCoMPaT200 [1] provides surface points with part labels for each object. To extend these labels to interior shape points, we query their $5$ nearest points and assign the label via a majority vote. Once part labels are obtained, we map corresponding material attributes based on part annotations in our material-augmented dataset.
B.4 Perception-based Baseline
In Section 5.2, we introduced the perception-based baseline (i.e., “B. Perc.”). The baseline uses perceptual models to estimate material properties based on an off-the-shelf 3D generation model. Here, we present more details of how this baseline is implemented. Specifically, we adopt TRELLIS [79], a state-of-the-art 3D generation model, to generate textured 3D shapes given image or text conditions. To obtain the material properties of each generated shape, we then leverage an open-vocabulary 3D part segmentation model, Find3D [45], to get the part labels for the sampled surface points. Note that the query part names we provide to Find3D are derived from the set of part labels in our dataset, which comes from 3DCoMPaT200 [1]. Finally, we leverage ChatGPT-4o [50] by providing it with two renderings of the generated 3D object and asking it to estimate the material properties for each part of the object retrieved by Find3D. The prompts we employed follow the one used in our dataset collection pipeline, as described in Section D.2. After we acquire the material information for each part of the 3D object, we propagate this information to the points sampled within the object’s volume using the same strategy outlined in Section B.3. Finally, we simulate the generated object with the estimated material properties.
B.5 Training Details
For object autoencoding, we use a surface point cloud of $2,048$ points as input to the autoencoder. At each iteration, we sample $1,024$ query points from the bounding sphere of the object and another $1,024$ points from the surface region for attribute prediction. Following [86], we ensure an equal distribution of query points between occupied and non-occupied regions. We set the weight coefficients described in Equation 6 as $\lambda_{c}=\lambda_{\nu}=\lambda_{\phi}=\lambda_{\rho}=\lambda_{M}=0.1$, $\lambda_{E}=\lambda_{\sigma}=0.01$, and $\lambda_{r}=0.001$. The autoencoder is trained with a batch size of $256$ for $500$ epochs, using a learning rate that linearly increases to $10^{-4}$ over the first $25$ epochs before gradually decaying to $10^{-5}$ following a cosine schedule.
For object generation, we train our diffusion models with a batch size of $256$ for $3,000$ epochs. The learning rate is linearly increased to $2\times 10^{-4}$ over the first $200$ epochs and then gradually decays to $10^{-6}$ using a cosine schedule. We adopt the default hyperparameter settings of EDM [26]. During sampling, the final latent codes are obtained using only 18 denoising steps. All our training experiments are conducted on two NVIDIA L40s GPUs.
Appendix C Dataset Details
Train 75 418 898 27 18 27 139 71 383 278 37 91 2462
Valid 10 26 52 5 2 5 21 11 21 18 6 3 180
Test 24 48 106 8 7 10 39 18 46 31 11 14 362
Table 4: Data composition of our proposed dataset.
Figure 10: More dataset statistics.
Figure 11: Statistics of fine-grained material types in our dataset.
In Table 4, we present the detailed composition of the proposed dataset, which covers $12$ common categories with part-level material annotations for $3,004$ shapes. After the VLM-guided material annotation, we obtained $8$, $7$, $10$ types of fine-grained materials for “fabric”, “leather”, and “plastic” respectively [^1] We did not obtain parts labeled with “bonded leather” for the “leather” type, and parts labeled with “flexible polyvinyl chloride”, “polystyrene” for the “plastic” type from the VLM.. The occurrence of parts labeled with these fine-grained materials is shown in Figure 11.
Furthermore, we present more dataset statistics in Figure 10. We begin by illustrating the number of rectification attempts made by experts for each object during the data verification phase outlined in Section 3.2. As presented in Figure 10(a), approximately one-third of the data has received at least one round of material parameter updates based on the experts’ feedback [^2] Please refer to Figure 13, where we provide some screenshots of the web interface we used for verifying material annotations.. This highlights the importance of human verification in correcting the initial proposals generated by the VLM. Furthermore, it is noteworthy that most verifications are completed within two iterations, which demonstrates the efficiency of our data annotation pipeline. Next, we provide the statistics of several material properties, including material behavior types in Figure 10(b), the logarithm of Young’s modulus in Figure 10(c), Poisson’s ratio in Figure 10(d), and the logarithm of yield stress in Figure 10(e). Specifically, M0 represents pure elastic materials using neo-Hookean elasticity and identity plasticity. M1 represents plastic materials with softening implemented with neo-Hookean elasticity and von Mises plasticity with damage. M2 is for plastic material without softening, which is supported by neo-Hookean elasticity and von Mises plasticity. Finally, M3 stands for granular material with StVK elasticity and Drucker-Prager plasticity as the material models. These figures showcase the material diversity of our collected dataset.
In summary, our dataset offers detailed material properties, which serve as a first step for advancements in simulation-driven learning, physics-informed generative modeling, and interactive virtual environments. We hope that our dataset can expand the scope of 3D perception research and contribute to developing effective algorithms and techniques that improve our understanding of the physical world.
Figure 12: Sampled simulation results from our proposed dataset.
Figure 13: Screenshots of the interactive panel for verifying material annotations.
Appendix D Full Prompts
@@
In this section, we provide the detailed prompts we used in the VLM-guided material annotation stage described in Section 3.
D.1 Fine-grained Material Types
@@
"part_name": "most_suitable_material_type",
@@
D.2 Material Models and Parameters
@@ 3. Drucker-Prager plasticity (Youngs modulus, Poissons ratio, friction angle);
For each part, you need to specify **both the elastic and the plastic** material model and the material parameters, such as Youngs modulus, Poissons ratio, density, etc.
Please provide your answer in the following JSON format (For Youngs modulus and yield stress, the unit is Pa. For density, the unit is kg/m^3.):
@@
"...": ..., // other parameters, e.g., yield stress ("sigma_y"), friction angle ("phi"), density ("rho")
@@
D.3 Update with Expert Feedback
In this case, we send three messages with the role set by user, assistant, user in succession. The first user prompt is the same as the prompt for obtaining the initial material properties. The assistant prompt is the output of the original query from the VLM. The second user prompt is:
@@ D.3 Update with Expert Feedback
Appendix E Limitations
Our current dataset of annotated shapes includes 3K objects and 15.5K labeled parts. This size is relatively small, especially for training large-scale 3D or 4D foundation models. We believe that our semi-automatic pipeline can be enhanced in the future to generate significantly larger datasets. Artifacts in 3D shape and texture generation are sometimes visible in our method (e.g., see the left, dark-shaded armrest of the first sofa shown in Figure 8). Generating alternative representations based on our physics-aware latents and denoiser, such as meshes [79] or marching tetrahedra [59], along with the generation of high-resolution textures and physically based rendering materials (PBR) [61], in an end-to-end manner could help reduce artifacts and further enhance the quality of the results.